# Greedy Heuristics for Sampling-based Motion Planning in High-Dimensional State Spaces

### v1

@@ 4.2 Greedy Informed Sampling
G-RRT* finds better solution paths to a planning problem by growing two trees, $G_{a}=(V_{a},E_{a})$ from the start and $G_{b}=(V_{b},E_{b})$ from the goal, consisting of the set of vertices, $V_{a}\cup V_{b}$, and edges, $E_{a}\cup E_{b}$, in free space towards randomly sampled states. These graphs are rewired with newly added vertices in such a way that the cost to come of the nearby vertices are minimized. Like Informed RRT*, once an initial solution is found, it only focuses the random sampling on the subset of the problem domain with admissible ellipsoidal heuristics, thereby increasing the convergence rate towards the optimum. However, instead of exploring all the states within the informed subset that can improve the current solution cost, G-RRT* differs from Informed RRT* in that it exploits the current solution path information and biases the random sampling on the decreasing subset of the hyperellipsoid (the $L^{2}$ greedy informed set). It also samples from the informed exploration subset to maintain the uniform sample distribution for asymptotic optimality guarantees. Specifically, G-RRT* balances the exploration and exploitation of the random sampling process by introducing a parameter, $\epsilon\in[0,1]$, called the greedy biasing ratio. If $\epsilon=0$, it only focuses the sampling on the informed subset, behaving the same as Informed RRT* algorithm with pure exploration. On the other hand, if $\epsilon=1$, the algorithm exploits the current solution path information and only finds the set of states within the greedy informed set with no exploration. Therefore, a good balance between uniform sampling for exploration and path-biased sampling for exploitation is necessary to rapidly reduce the size of the informed subset for faster convergence rate and global optimality guarantees.
@@
At each iteration, G-RRT* finds the traverse diameter of the respective hyperellipsoids for random sampling process (Alg. 1, Line 1). First, it finds the current best goal vertex, $\mathbf{x}_{\text{i}}$, with the minimum cost-to-come value from the set of vertices in the goal region (Alg. 2, Line 2). With probability $\epsilon$, the algorithm determines the current maximum cost, $c_{\text{i}}$, for the sampling process by greedily selecting the best vertex in the existing solution path with maximum admissible heuristics (i.e., exploitation with $L^{2}$ greedy informed set) (Alg. 2, Lines 22). With probability $1-\epsilon$, it defines the current maximum cost by determining the cost required to reach the current best goal vertex (i.e., exploration with $L^{2}$ informed set) (Alg. 2, Lines 2 and 2). The uniformly distributed random state can then be sampled from the hyperellipsoid if there exists a solution path (Alg. 1, Line 1, Alg. 3, Lines 33) or from the problem state space if no solution is found yet (Alg. 1, Line 1, Alg. 3, Lines 33).
(a) Parent selection process of G-RRT*
(b) Rewiring process of G-RRT*
Figure 4: Incremental Q-RRT*-style greedy parent selection and tree rewiring processes of G-RRT* algorithm. The local neighbourhood of the newly added vertex $\mathbf{x}_{\text{new}}$ is highlighted in circle with a dashed line while nearest vertices as well as their ancestors are denoted as gray dots. Black dots represent the rest of the vertices in the tree which does not include within a certain depth level.
4.3 Tree Extension
@@
Once a new random sample has been found, G-RRT* performs the tree extension process just as in RRT-Connect by connecting one tree towards the other and alternating between the two trees repeatedly. First, the selected tree is extended from the nearest vertex $\mathbf{x}_{\text{nearest}}$ towards $\mathbf{x}_{\text{rand}}$ using the Steer function also known as a local planner, resulting in $\mathbf{x}_{\text{new}}$ (Alg. 1, Line 1, Alg. 4, Lines 44). If the resulting edge from the nearest vertex $\mathbf{x}_{\text{nearest}}$ to the new vertex $\mathbf{x}_{\text{new}}$ is valid, $\mathbf{x}_{\text{new}}$ is added to the set of vertices $V$ of the current tree (Alg. 1, Line 1, Alg. 4, Lines 44). To achieve almost-sure asymptotic optimality, G-RRT* does the incremental tree rewiring techniques in the local neighbour hood of the newly added vertex $\mathbf{x}_{\text{new}}$ as in RRT*. However, G-RRT* also considers outside of the neighbour hood, i.e., it takes into consideration of the ancestors of $X_{\text{near}}$ up to a certain depth level similar to Q-RRT* algorithm. This parent selection process tends to result in more straighten tree edges with lower cost initial solutions compared to only using the nearby vertices $X_{\text{near}}$. Similarly, these ancestors are also considered during the tree rewiring procedure to check whether they can improve any of $\mathbf{x}_{\text{new}}$’s neighbours. Figure 4 illustrates the greedy parent selection and rewiring processes of the G-RRT* algorithm.
Definition 8 ((Predecessor of a vertex)).
Let $\mathbf{v}$ be a vertex in a tree $G=(V,E)$ defined by a set of vertices $V$ and edges $E$. We then define the predecessor of $\mathbf{v}$, denoted by $\text{pred}({\mathbf{v})}=\mathbf{u}$, as a parent vertex of $\mathbf{v}$ which has a directed edge leading into $\mathbf{v}$ in $G$ where $\mathbf{u},\mathbf{v}\in V$ and $(\mathbf{u},\mathbf{v})\in E$.
Definition 9 ((Ancestor of a vertex)).
Let $\mathbf{v}$ be a vertex in a tree $G=(V,E)$ defined by a set of vertices $V$ and edges $E$. Let $d$ be the depth level above $\mathbf{v}$ in the tree. We then define the ancestor of a vertex at $d$-th parent as the vertex that is $d$ levels above $\mathbf{v}$ in the tree, following the path towards the root start vertex $\mathbf{x}_{I}$:
$\text{ancestor}_{d}(\mathbf{v})=\text{pred}^{(d)}(\mathbf{v})$ (12)
where $\text{pred}^{(d)}(\mathbf{v})$ represents recursively applying the predecessor function $d$ times starting from vertex $\mathbf{v}$.
Definition 10 ((Ancestry set)).
Let $S$ be the set of vertices in a tree $G=(V,E)$ defined by a set of vertices $V$ and edges $E$. We define the ancestry set of $S$, denoted by $X_{\text{ancestors}}(S)$, as the set containing all vertices that are ancestors of any vertex in $S$ up to the $d$-th parent. Formally, for $d\geq 1$, the ancestry set is given by:
$X_{\text{ancestors}}(S)=\bigcup_{\mathbf{v}\in S}\bigcup_{i=1}^{d}\text{ancestor}_{i}(\mathbf{v})$ (13)
Once a new vertex $\mathbf{x}_{\text{new}}$ is added into the tree, G-RRT* finds the best potential parent vertex for this newly added state by considering the cost-to-come to a set of nearby vertices defined by the local neighborhood radius (Alg. 1, Line 1, Alg. 4, Line 4). It also considers the ancestry set $X_{\text{ancestors}}$ of $X_{\text{near}}$ (Definition 10) up to a certain depth level during this parent selection process (Alg. 1, Line 1, Alg. 4, Line 4). If connecting to a vertex from these candidate parent sets improves the current cost-to-come to a newly added vertex $\mathbf{x}_{\text{new}}$ and the edge is valid, this resulting edge from the best parent vertex $\mathbf{x}_{\text{min}}$ to $\mathbf{x}_{\text{new}}$ is added to a set of edges $E$ of the current tree (Alg. 1, Line 1, Alg. 4, Lines 44).
@@
G-RRT* also considers this ancestry set $X_{\text{ancestors}}$ during its rewiring process. If connecting to either a new vertex $\mathbf{x}_{\text{new}}$ or one of the ancestors can lower the cost-to-come of the nearby vertices, they are rewired to be the descendants of this vertex if the new edge is collision-free (Alg. 1, Line 1, Alg. 4, Lines 44). Similar to RRT-Connect algorithm, tree extension process of G-RRT* also results in three output states: REACHED if a newly generated vertex $\mathbf{x}_{\text{new}}$ reaches an input vertex $\mathbf{x}$ (Alg. 1, Line 1, Alg. 4, Lines 44), ADVANCED if $\mathbf{x}_{\text{new}}\neq\mathbf{x}$ is added to the vertex set (Alg. 1, Line 1, Alg. 4, Lines 44) and TRAPPED if the edge from $\mathbf{x}_{\text{nearest}}$ to $\mathbf{x}_{\text{new}}$ is in collision (Alg. 1, Line 1, Alg. 4, Line 4).
4.4 Greedy Connect Heuristic
@@
Certain planning problems are extremely challenging for planners to find the paths especially when either start or goal states (or both) are blocked by enclosures or narrow passages (for example, problems in Figures 6(b) and 6(c)). This difficulty increases as the state dimension grows, making RRTs impractical to be used in higher state space applications. In our work, we tackle this problem by relying on a bi-directional search, maintaining two separate RRT trees at the start and goal states respectively. In each iteration, while one tree is extended toward the randomly generated state using the EXTEND* procedure (Section 4.3), G-RRT* attempts to connect the newly added vertex to the nearby vertex of the other tree using the greedy CONNECT heuristic similar to RRT-Connect algorithm. It then alternates between the steps by swapping the two trees. This approach helps G-RRT* to find paths quickly in these challenging situations by using the benefits of uniform explorative nature of RRTs to avoid the local minima pitfalls while exploiting with greedy connection strategy to connect the two trees.
@@
Once a tree extension process towards the randomly generated vertex is succeeded, G-RRT* attempts to extend the other tree towards the newly added vertex of the current tree using the greedy CONNECT* heuristic (Alg. 1, Lines 11). This process iteratively calls EXTEND* until the other tree either reaches a newly added vertex or fails to do the extension anymore due to the collision in edge (Alg. 5). If this connection is succeeded, i.e., the other tree successfully connects the newly added vertex of the current tree, solution path is found and this edge is added to the solution edge set $E_{\text{sol'n}}$ which maintains all the edges connecting the two trees (Alg. 1, Lines 11). In order to make sure both trees are growing towards each other greedily, G-RRT* swaps the two trees before proceeding the next iteration of the algorithm (Alg. 1, Line 1).
5 Analysis
Figure 5: Illustration of the sub-optimality of greedy informed set in an example planning scenario. The $L^{2}$ greedy informed set, $\mathcal{X}_{\hat{f}_{\text{greedy}}}$, highlighted in a dashed ellipse, is constructed based on the current tortuous solution path, $\pi_{i}$. The ture optimal path, $\pi^{*}$, is depicted as a dotted line. Here, the greedy informed set does not contain all the vertices that can improve the current solution cost (i.e., those that contribute to approaching $\pi^{*}$ and lie outside of the hyperellipsoid) due to the nature of its greedy exploitation.
5.1 Probabilistic completeness
@@
A path planning algorithm is considered robustly feasible if there exists a solution path with strong $\delta$-clearance (i.e., a distance $\delta$ away from any obstacle in $\mathcal{X}\in\mathbb{R}^{n}$) for some $\delta>0$ (Karaman and Frazzoli, 2011). An algorithm for solving any robustly feasible planning problem ($\mathcal{X}_{\text{free}},\mathbf{x}_{I},\mathbf{x}_{G}$) has probabilistic completeness guarantees if there exists a goal vertex $\mathbf{x}_{G}$ in the graph where a start vertex $\mathbf{x}_{I}$ becomes connected to $\mathbf{x}_{G}$ as the number of samples approaches infinity (Definition 3). In other words, the algorithm is said to be probabilistically complete, if given sufficient time, it can find a solution to the path planning problem if one exists.
@@
5.2 Asymptotic optimality
Consider a robustly feasible path planning problem defined by the state space $\mathcal{X}_{\text{free}}$, initial state $\mathbf{x}_{I}$, and goal state $\mathbf{x}_{G}$, we define the optimal path planning problem as to find a path $\pi^{*}$ from $\mathbf{x}_{I}$ to $\mathbf{x}_{G}$ that minimizes a cost function $c:\Sigma\rightarrow\mathbb{R}^{+}_{0}$ where $\Sigma$ is the set of all feasible paths (Definition 2). A solution path $\pi^{*}$ is considered robustly optimal if there exists another path $\pi_{0}$ with strong $\delta$-clearance in the same homotopy class where $c(\pi_{0})=\min\{c(\pi):\pi\text{ is feasible}\}$ (Karaman and Frazzoli, 2011). A solution path $\pi^{*}$ is considered asymptotically optimal if, for any robustly feasible planning problem ($\mathcal{X}_{\text{free}},\mathbf{x}_{I},\mathbf{x}_{G}$) and cost function $c:\Sigma\rightarrow\mathbb{R}^{+}_{0}$ that produces a robustly optimal solution, the following holds:
@@ 5.2 Asymptotic optimality
RRT* has proven to be asymptotically optimal by incrementally rewiring the tree around new vertices (Karaman and Frazzoli, 2011). These are the set of vertices which are within the $n$-dimensional unit ball of radius, $r_{rewire}$, of the newly added vertex in the Euclidean space:
@@
Direct informed sampling approaches have also shown to be maintaining the property of asymptotic optimality of RRT* however they only search on a subset of the original planning problem (Gammell et al., 2014, 2018). Therefore, this reduces the computational requirements of rewiring at each planning iteration due to the smaller number of rewiring neighborhoods. Here, the rewiring requirements to maintain the asymptotic optimality in the smaller subset of the planning domain is a function of the number of vertices in the informed set, $|V\cap\mathcal{X}_{\hat{f}}|$, and its measure, $\lambda(\mathcal{X}_{\hat{f}})$.
G-RRT* uses the greedy informed set $\mathcal{X}_{\hat{f}_{\text{greedy}}}$ described in Section 3 to focus the search on even smaller subset of the planning domain than the informed set, where, it is bounded from above by the informed subset, i.e., $\mathcal{X}_{\hat{f}_{\text{greedy}}}\supseteq\mathcal{X}_{\hat{f}}$. This increases the chances of finding the vertices that are more likely to be inside the omniscient set (Definition 5), improving the rate of converging to the optimal solution. Here, the rewiring requirements of G-RRT* is a function of the number of vertices in the greedy informed set, $|V\cap\mathcal{X}_{\hat{f}_{\text{greedy}}}|$, and its measure, $\lambda(\mathcal{X}_{\hat{f}_{\text{greedy}}})$ where
@@
However, not all vertices in the greedy informed set is guaranteed to be in the omniscient set since it is an overly approximated version of it which may or may not be an upper bound estimate of the true omniscient set (see Figure 5). In other words, vertices that can lead to better solution paths may exist in outside homotopy regions of the greedy informed set. Since G-RRT* also samples from the informed subset based on the greedy biasing ratio (See Section 4.2), it still holds the asymptotic optimality guarantees of direct informed sampling techniques.
@@
Figure 7: Planner performances versus running time on the many homotopy classes problem described in Section 6.1 (Figure 6(a)). The success plots represent the percentage of number of successful runs over time and the cost plots show the median solution cost versus run time for each planner. The error bars and shaded areas denote non-parametric 99% confidence intervals on the median. Each planner was run for $100$ trials with running times of $0.5$, $5.0$, $30.0$ and $100.0$ seconds in $\mathbb{R}^{2}$, $\mathbb{R}^{4}$, $\mathbb{R}^{8}$ and $\mathbb{R}^{16}$, presented in plots (a), (b), (c) and (d), respectively.
@@
Figure 8: Planner performances versus running time on the narrow passage gap problem described in Section 6.1 (Figure 6(b)). The success plots represent the percentage of number of successful runs over time and the cost plots show the median solution cost versus run time for each planner. The error bars and shaded areas denote non-parametric 99% confidence intervals on the median. Each planner was run for $100$ trials with running times of $0.5$, $1.0$, $5.0$ and $30.0$ seconds in $\mathbb{R}^{2}$, $\mathbb{R}^{4}$, $\mathbb{R}^{8}$ and $\mathbb{R}^{16}$, presented in plots (a), (b), (c) and (d), respectively.
@@
Figure 9: Planner performances versus running time on the double enclosure problem described in Section 6.1 (Figure 6(c)). The success plots represent the percentage of number of successful runs over time and the cost plots show the median solution cost versus run time for each planner. The error bars and shaded areas denote non-parametric 99% confidence intervals on the median. Each planner was run for $50$ trials with running times of $10.0$ and $300.0$ seconds in $\mathbb{R}^{4}$ and $\mathbb{R}^{8}$, presented in plots (a) and (b), respectively.
@@
Figure 11: Planner performances versus running time on the manipulation experiment described in Section 6.2 (Figure 10). The success plots represent the percentage of a number of successful runs over time and the cost plots show the median solution cost versus run time for each planner. The error bars and shaded areas denote non-parametric 99% confidence intervals on the median. Each planner was run for $50$ trials with running time of $100.0$ seconds in $\mathbb{R}^{14}$.
6 Simulation and Experiments

### v2

@@ IV-B Greedy Informed Sampling
G-RRT* finds better solution paths to a planning problem by growing two trees, $G_{a}=(V_{a},E_{a})$ from the start and $G_{b}=(V_{b},E_{b})$ from the goal, consisting of the vertices $V_{a}\cup V_{b}$ and edges $E_{a}\cup E_{b}$, towards randomly sampled states in free space. These graphs are rewired with newly added vertices to minimize the cost-to-come for nearby vertices. Similar to Informed RRT*, once an initial solution is found, the algorithm focuses random sampling on a subset of the problem domain, constrained by admissible ellipsoidal heuristics, which accelerates convergence to the optimal solution. However, unlike Informed RRT*, which explores all states in the informed subset to improve the current solution cost, G-RRT* leverages information from the current solution path and biases random sampling toward a shrinking subset of the hyperellipsoid, known as the $L^{2}$ greedy informed set. It also samples from the informed exploration subset to maintain uniform distribution, ensuring asymptotic optimality. Specifically, G-RRT* balances the exploration and exploitation of the random sampling process by introducing a parameter, $\epsilon\in[0,1]$, called the greedy biasing ratio. If $\epsilon=0$, it only focuses the sampling on the informed subset, behaving the same as the Informed RRT* algorithm with pure exploration. On the other hand, if $\epsilon=1$, the algorithm exploits the current solution path information and only finds the set of states within the greedy informed set with no exploration. Therefore, finding a good balance between uniform sampling for exploration and path-biased sampling for exploitation is necessary to rapidly reduce the size of the informed subset for faster convergence rates and global optimality guarantees.
@@
At each iteration, G-RRT* finds the transverse diameter of the corresponding hyperellipsoids for the random sampling process (Alg. 1, Line 1; Alg. 2). First, it determines the current best cost, $c_{i}$, by calculating the cost required to reach the best goal vertex from the set of vertices in the goal region (Alg. 2, Lines 22). With probability $\epsilon$, the algorithm computes the current maximum cost, $c_{\text{max}}$, for the sampling process by greedily selecting the best vertex along the existing solution path with the maximum admissible heuristic cost (i.e., exploitation using the $L^{2}$ greedy informed set) (Alg. 2, Lines 22). This greedy maximum cost is recalculated only when a new, better solution is found (Alg. 2, Line 2) to improve performance. With probability $1-\epsilon$, the current best cost is defined by the solution cost computed in Alg. 2, Line 2 (i.e., exploration using the $L^{2}$ informed set), and the current minimum solution cost is cached for the next iteration (Alg. 2, Lines 2 and 2).
IV-C Tree Extension
@@
Once a new random sample is found, G-RRT* performs the tree extension process similar to RRT-Connect by alternating between the two trees and connecting one tree toward the other repeatedly. First, the selected tree is extended from the nearest vertex, $\mathbf{x}_{\text{nearest}}$, toward $\mathbf{x}_{\text{rand}}$ using the Steer function, also known as a local planner, resulting in $\mathbf{x}_{\text{new}}$ (Alg. 1, Line 1; Alg. 4, Lines 4 and 4). If the resulting edge from $\mathbf{x}_{\text{nearest}}$ to $\mathbf{x}_{\text{new}}$ is valid, $\mathbf{x}_{\text{new}}$ is added to the set of vertices, $V$, of the current tree (Alg. 1, Line 1; Alg. 4, Lines 4 and 4). To achieve almost-sure asymptotic optimality, G-RRT* performs incremental tree rewiring in the local neighborhood of the newly added vertex, $\mathbf{x}_{\text{new}}$, as in RRT*. Once $\mathbf{x}_{\text{new}}$ is added to the tree, G-RRT* searches for the best potential parent vertex for this newly added state by considering the cost-to-come to nearby vertices within the local neighborhood radius (Alg. 1, Line 1; Alg. 4, Line 4). If connecting to a vertex from this set of candidate parents reduces the current cost-to-come to $\mathbf{x}_{\text{new}}$ and the edge is valid, the edge from the best parent vertex, $\mathbf{x}_{\text{min}}$, to $\mathbf{x}_{\text{new}}$ is added to the set of edges, $E$, in the current tree (Alg. 1, Line 1; Alg. 4, Lines 44).
@@
If connecting to the new vertex, $\mathbf{x}_{\text{new}}$, can lower the cost-to-come of nearby vertices, they are rewired to become descendants of this vertex, provided the new edge is collision-free (Alg. 1, Line 1; Alg. 4, Lines 44). Similar to the RRT-Connect algorithm, the tree extension process in G-RRT* results in three possible outcomes: REACHED, if the newly generated vertex, $\mathbf{x}_{\text{new}}$, reaches the input vertex $\mathbf{x}$ (Alg. 1, Line 1; Alg. 4, Lines 4 and 4), ADVANCED, if $\mathbf{x}_{\text{new}}\neq\mathbf{x}$ is added to the vertex set (Alg. 1, Line 1; Alg. 4, Lines 4 and 4) and TRAPPED, if the edge from $\mathbf{x}_{\text{nearest}}$ to $\mathbf{x}_{\text{new}}$ is in collision (Alg. 1, Line 1; Alg. 4, Line 4).
IV-D Greedy Connect Heuristic
@@
Certain planning problems are extremely challenging for planners to solve, particularly when either the start or goal states (or both) are obstructed by enclosures or narrow passages (as in Figures 7(b) and 7(c)). This difficulty increases as the dimensionality of the state space grows, making RRTs impractical for higher-dimensional applications. In our work, we address this challenge by utilizing a bi-directional search, maintaining two separate RRT trees at the start and goal states, respectively. In each iteration, while one tree is extended toward a randomly generated state using the EXTEND* procedure (Section IV-C), G-RRT* attempts to connect the newly added vertex to a nearby vertex from the other tree using a greedy CONNECT heuristic, similar to the RRT-Connect algorithm. It then alternates between these steps by swapping the two trees. This approach enables G-RRT* to quickly find paths in these challenging situations by leveraging the uniform explorative nature of RRTs to avoid local minima pitfalls while also exploiting a greedy strategy to connect the two trees.
@@
Once the tree extension process toward the randomly generated vertex succeeds, G-RRT* attempts to extend the other tree toward the newly added vertex of the current tree using the greedy CONNECT* heuristic (Alg. 1, Lines 11). This process iteratively calls EXTEND* until the other tree either reaches the newly added vertex or fails to extend further due to a collision along the edge (Alg. 5). If this connection succeeds, meaning the other tree successfully connects to the newly added vertex of the current tree, a solution path is found, and this edge is added to the solution edge set $E_{\text{sol'n}}$, which maintains all the edges connecting the two trees (Alg. 1, Lines 1 and 1). To ensure both trees grow toward each other greedily, G-RRT* swaps the two trees before proceeding to the next iteration of the algorithm (Alg. 1, Line 1).
V Analysis
V-A Probabilistic completeness
@@
A path planning algorithm is considered robustly feasible if there exists a solution path with strong $\delta$-clearance (i.e., a distance $\delta$ away from any obstacle in $\mathcal{X}\in\mathbb{R}^{n}$) for some $\delta>0$ [8]. An algorithm for solving any robustly feasible planning problem ($\mathcal{X}_{\text{free}},\mathbf{x}_{I},\mathbf{x}_{G}$) has probabilistic completeness guarantees if there exists a goal vertex $\mathbf{x}_{G}$ in the graph where a start vertex $\mathbf{x}_{I}$ becomes connected to $\mathbf{x}_{G}$ as the number of samples approaches infinity (Definition 3). In other words, the algorithm is said to be probabilistically complete, if given sufficient time, it can find a solution to the path planning problem if one exists.
@@
V-B Asymptotic optimality
Consider a robustly feasible path planning problem defined by the state space $\mathcal{X}_{\text{free}}$, initial state $\mathbf{x}_{I}$, and goal state $\mathbf{x}_{G}$, we define the optimal path planning problem as to find a path $\pi^{*}$ from $\mathbf{x}_{I}$ to $\mathbf{x}_{G}$ that minimizes a cost function $c:\Sigma\rightarrow\mathbb{R}_{\geq 0}$ where $\Sigma$ is the set of all feasible paths (Definition 2). A solution path $\pi^{*}$ is considered robustly optimal if there exists another path $\pi_{0}$ with strong $\delta$-clearance in the same homotopy class where $c(\pi_{0})=\min\{c(\pi):\pi\text{ is feasible}\}$ [8]. A solution path $\pi^{*}$ is considered asymptotically optimal if, for any robustly feasible planning problem ($\mathcal{X}_{\text{free}},\mathbf{x}_{I},\mathbf{x}_{G}$) and cost function $c:\Sigma\rightarrow\mathbb{R}_{\geq 0}$ that produces a robustly optimal solution, the following holds:
@@ V-B Asymptotic optimality
RRT* has proven to be asymptotically optimal by incrementally rewiring the tree around new vertices [8]. These are the set of vertices which are within the $n$-dimensional unit ball of radius, $r_{rewire}$, of the newly added vertex in the Euclidean space:
@@
Direct informed sampling approaches have also shown to be maintaining the property of asymptotic optimality of RRT* however they only search on a subset of the original planning problem [12, 31]. Therefore, this reduces the computational requirements of rewiring at each planning iteration due to the smaller number of rewiring neighborhoods. Here, the rewiring requirements to maintain the asymptotic optimality in the smaller subset of the planning domain is a function of the number of vertices in the informed set, $|V\cap\mathcal{X}_{\hat{f}}|$, and its measure, $\lambda(\mathcal{X}_{\hat{f}})$.
G-RRT* uses the greedy informed set $\mathcal{X}_{\hat{f}_{\text{greedy}}}$ described in Section III to focus the search on even smaller subset of the planning domain than the informed set, where, it is bounded from above by the informed subset, i.e., $\mathcal{X}_{\hat{f}_{\text{greedy}}}\supseteq\mathcal{X}_{\hat{f}}$. This increases the chances of finding the vertices that are more likely to be inside the omniscient set (Definition 5), improving the rate of converging to the optimal solution. Here, the rewiring requirements of G-RRT* is a function of the number of vertices in the greedy informed set, $|V\cap\mathcal{X}_{\hat{f}_{\text{greedy}}}|$, and its measure, $\lambda(\mathcal{X}_{\hat{f}_{\text{greedy}}})$ where
@@
However, not all vertices in the greedy informed set are guaranteed to be in the omniscient set as it is an overly approximated version of it which may or may not be an upper bound estimate of the true omniscient set (see Figure 6). In other words, vertices that could lead to better solution paths may exist outside the homotopy regions of the greedy informed set (Theorem 2). Since G-RRT* also samples from the informed subset based on the greedy biasing ratio (Section IV-B), it still holds the asymptotic optimality guarantees of direct informed sampling techniques.
This raises the question of how sampling with the greedy informed set affects the worst-case number of samples required for probabilistic optimality. Theorem 3 quantifies this impact by comparing it to sampling with the informed set only.
Theorem 3 (Impact on the worst-case number of samples for probabilistic optimality).
The worst-case number of samples required for probabilistic optimality using the greedy informed set grows proportionally by a factor of $\frac{1}{1-\epsilon}$ compared to using only the informed set.
Proof.
Let $P(\mathbf{x}_{\text{rand}}\in\mathcal{X}_{f})$ represent the probability of sampling a state from the omniscient set. According to Lemma 5 from [31], sampling states from the omniscient set is a necessary condition for improving the current solution in RRT*-like algorithms. Let $P(\mathbf{x}_{\text{rand}}\in\mathcal{X}_{\hat{f}})$ represent the probability of sampling a state from the informed set. Since $\mathcal{X}_{\hat{f}}\supseteq\mathcal{X}_{f}$, as noted in [31], sampling from the informed set is also a necessary condition for probabilistic optimality, and its probability is bounded below by the probability of sampling from $\mathcal{X}_{f}$:
$P(\mathbf{x}_{\text{rand}}\in\mathcal{X}_{f})\leq P(\mathbf{x}_{\text{rand}}\in\mathcal{X}_{\hat{f}})$
Thus, the expected number of samples required to improve the current solution to a holonomic problem depends on the probability of sampling the informed set, and is given by:
$E[n]=\frac{1}{P(\mathbf{x}_{\text{rand}}\in\mathcal{X}_{\hat{f}})}$ (18)
Recall that with probability $1-\epsilon$, we sample from the informed set, and with probability $\epsilon$, we sample from the greedy informed set. Therefore, the overall probability of sampling from $\mathcal{X}_{\hat{f}}$ in any given iteration is reduced, since we do not always sample from $\mathcal{X}_{\hat{f}}$. Specifically, the new probability $P_{\text{new}}(\mathbf{x}_{\text{rand}}\in\mathcal{X}_{\hat{f}})$ of selecting a random sample from $\mathcal{X}_{\hat{f}}$ is:
$P_{\text{new}}(\mathbf{x}_{\text{rand}}\in\mathcal{X}_{\hat{f}})=(1-\epsilon)\cdot P(\mathbf{x}_{\text{rand}}\in\mathcal{X}_{\hat{f}})$ (19)
Substituting (19) into (18), the expected number of samples required to sample from $\mathcal{X}_{\hat{f}}$ now becomes:
$E[n_{\text{new}}]$ $=\frac{1}{P_{\text{new}}(\mathbf{x}_{\text{rand}}\in\mathcal{X}_{\hat{f}})}$ (20)
$=\frac{1}{(1-\epsilon)\cdot P(\mathbf{x}_{\text{rand}}\in\mathcal{X}_{\hat{f}})}$
Comparing (18) and (20):
$\frac{E[n_{\text{new}}]}{E[n]}=\frac{\frac{1}{(1-\epsilon)\cdot P(\mathbf{x}_{\text{rand}}\in\mathcal{X}_{\hat{f}})}}{\frac{1}{P(\mathbf{x}_{\text{rand}}\in\mathcal{X}_{\hat{f}})}}=\frac{1}{1-\epsilon}$
Therefore, in the worst case where the greedy informed set does not include states from the optimal solution, the number of samples required for asymptotic optimality increases by a factor of $\frac{1}{1-\epsilon}$ compared to only sampling from $\mathcal{X}_{\hat{f}}$. ∎
@@
Figure 8: Planner performance versus running time on the many homotopy classes problem described in Section VI-A (Figure 7(a)). The success plots represent the percentage of number of successful runs over time and the cost plots show the median solution cost versus run time for each planner. The error bars and shaded areas denote non-parametric 99% confidence intervals on the median. Each planner was run for $100$ trials with running times of $0.5$, $5.0$, $30.0$ and $100.0$ seconds in $\mathbb{R}^{2}$, $\mathbb{R}^{4}$, $\mathbb{R}^{8}$ and $\mathbb{R}^{16}$, presented in plots (a), (b), (c) and (d), respectively.
@@
Figure 9: Planner performance versus running time on the narrow passage gap problem described in Section VI-A (Figure 7(b)). The success plots represent the percentage of number of successful runs over time and the cost plots show the median solution cost versus run time for each planner. The error bars and shaded areas denote non-parametric 99% confidence intervals on the median. Each planner was run for $100$ trials with running times of $0.5$, $1.0$, $5.0$ and $30.0$ seconds in $\mathbb{R}^{2}$, $\mathbb{R}^{4}$, $\mathbb{R}^{8}$ and $\mathbb{R}^{16}$, presented in plots (a), (b), (c) and (d), respectively.
@@
Figure 10: Planner performance versus running time on the double enclosure problem described in Section VI-A (Figure 7(c)). The success plots represent the percentage of number of successful runs over time and the cost plots show the median solution cost versus run time for each planner. The error bars and shaded areas denote non-parametric 99% confidence intervals on the median. Each planner was run for $50$ trials with running times of $10.0$ and $300.0$ seconds in $\mathbb{R}^{4}$ and $\mathbb{R}^{8}$, presented in plots (a) and (b), respectively.
@@
Figure 12: Planner performance versus running time on the manipulation experiment described in Section VI-B (Figure 11). The success plots represent the percentage of a number of successful runs over time and the cost plots show the median solution cost versus run time for each planner. The error bars and shaded areas denote non-parametric 99% confidence intervals on the median. Each planner was run for $50$ trials with running time of $100.0$ seconds in $\mathbb{R}^{14}$.
VI Simulation and Experiments